OpenStudy (anonymous):

How do I find functions of f(x) and g(x)? I'm currently struggling to understand how to do these problems for my Algebra 2 class. If anyone would be willing to walk me through the steps and explain why/how I do them that would be great! The functions f(x) and g(x) are shown below: f(x)=x^2-7 g(x)=x^2+3x+7 Show the algebraic processes explaining the steps used to determine the simplified expressions containing these functions as specified below. Part A: Find f(x)-g(x) Part B:A function, h(x) exisys such that h(x)=f(x)*g(x), What is the equation that function h(x) represents? Part C: Find f(x)

jimthompson5910 (jim_thompson5910):

f(x)-g(x) means you subtract ALL of g(x) from f(x)

jimthompson5910 (jim_thompson5910):

f(x)-g(x) = [ f(x) ] - [ g(x) ] f(x)-g(x) = [ x^2-7 ] - [ x^2+3x+7 ] f(x)-g(x) = x^2-7 - x^2-3x-7 ... distribute f(x)-g(x) = -3x - 14 ... combine like terms does that make sense?

OpenStudy (anonymous):

Yes it does! Can you explain how do I find part b? I don't get how to find the function of h(x)

jimthompson5910 (jim_thompson5910):

you will be multiplying f and g (instead of subtracting) what do you get when you do so?

OpenStudy (anonymous):

Would I get -3x *28?

jimthompson5910 (jim_thompson5910):

you will have this as part of your steps h(x) = f(x) * g(x) h(x) = [ f(x) ] * [ g(x) ] h(x) = [ x^2-7 ] * [ x^2+3x+7 ] what comes next?

OpenStudy (anonymous):

You would distribute, I got h(x)=x^4+3x^3-21x-49, is that right?

jimthompson5910 (jim_thompson5910):

that is correct

jimthompson5910 (jim_thompson5910):

"Part C: Find f(x)" makes no sense to me because they already gave you f(x)

OpenStudy (anonymous):

Oh I didn't finish typing part c, my bad! Part c: find f(x)+g(x)

jimthompson5910 (jim_thompson5910):

so you just add the functions now

jimthompson5910 (jim_thompson5910):

follow the same steps for subtracting, but replace the subtraction symbol with a plus symbol

OpenStudy (anonymous):

Is the answer for part c 2x^2+3x ?

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

Thank you so much! (:

jimthompson5910 (jim_thompson5910):

you're welcome